It is not obvious what is working.
It should be. As I mentioned, you can only trust the Stochmon signals between 0840 and 1130 UTC; before this time, the temperature controller was not connected, and after this time, c1sus is shut down and the MC is not locked, as you can see in your DV plot.
Within this time frame, the Stochmon signals are relatively stabilized (though there is some residual common-mode-ish drift since we are not using RFAM as the error signal---i.e., other things like laser power and frequency can mix in). Also, anytime after 0840, the controller signals behave as they should (these are unaffected by the status of c1sus):
I am concerned with the other issues that I mentioned in my previous post, namely: